The sidecap package (package op-

tion margincaption)

[figure]

Fig. 1: The dazed brown fox quickly gave 12345–67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

text area (odd page) margin area

Fig. 2: The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave

12345–67890 jumps!

[figure]

margin area text area (even page)